Ms Ananaya wanted to have a retirement corpus of Rs.1.5 Crores. She has 18 years to go
for retirement. But now she wants to retire 3 years earlier. What is the additional amount
that she needs to invest each year if the rate of return is 8%?
1
Expert's answer
2020-12-03T10:48:51-0500

this is an annuity

in the first case 18 years before retirement the payment is annual:

"A=\frac{15 000 000*0.08}{(1+0.08)^{18}-1)}=400 531.44"

in the second case, the payment will be 15 years before retirement:

"A=\frac{15 000 000*0.08}{(1+0.08)^{15}-1)}=552 443.17"

 additional amount:

552 443.17-400 531.44=151 911.73